Mullineaux Corporation has a target capital structure of 55 percent common stock and 45 percent debt. Its cost of equity is 11.1 percent, and the cost of debt is 5.8 percent. The relevant tax rate is 21 percent. What is the company’s WACC?
Answer #1

Answer:

After cost of debt = Cost of Debt * (1- tax rate)
After cost of Debt = 0.058 * (1 – 0.21)
After Cost of Debt = 0.058 * 0.79
After Cost of Debt = 0.04582 or 4.582%

WACC = (Weight of Common Stock * Cost of Equity) + (Weight of Debt * Cost of Debt)
WACC = (0.55 * 0.1110) + (0.45 * 0.04582)
WACC = 0.06105 + 0.020619
WACC = 0.08167
or WACC = 8.17%